I just help with part D. I am having a tough time understanding what I need to do, if you could explain it clearly that would be much appreciated! Thank you!(Measuring growth) Solarpower Systems earned $20 per share at the beginning of the year and paid out $9 in dividends to share
Answer #1

g=RoE*(1-payout rate)
Price=D0*(1+g)/(r-g)

1.
=21%*11/20=0.1155

2.
=9*(1+21%*11/20)/(14%-21%*11/20)=409.775510204082

3.
=13*(1+21%*7/20)/(14%-21%*7/20)=209.857142857143

4.
=2*(1+21%*18/20)/(14%-21%*18/20)=-48.530612244898